Brandon W.
asked 06/09/20
3 plain pencils and 5 pens in first box. 4 colored pencils and crayons in second box. One item from each box selected at random. What is probability that a pen and crayon are selected
Inactive Tutor
answered 06/09/20
5/8 * 4/8 = 5/8*1/2 = 5/16
Still looking for help? Get the right answer, fast.
OR
Find an Online Tutor Now
Choose an expert and meet online.
No packages or subscriptions, pay only for the time you need.
¢
€
£
¥
‰
µ
·
•
§
¶
ß
‹
›
«
»
<
>
≤
≥
–
—
¯
‾
¤
¦
¨
¡
¿
ˆ
˜
°
−
±
÷
⁄
×
ƒ
∫
∑
∞
√
∼
≅
≈
≠
≡
∈
∉
∋
∏
∧
∨
¬
∩
∪
∂
∀
∃
∅
∇
∗
∝
∠
´
¸
ª
º
†
‡
À
Á
Â
Ã
Ä
Å
Æ
Ç
È
É
Ê
Ë
Ì
Í
Î
Ï
Ð
Ñ
Ò
Ó
Ô
Õ
Ö
Ø
Œ
Š
Ù
Ú
Û
Ü
Ý
Ÿ
Þ
à
á
â
ã
ä
å
æ
ç
è
é
ê
ë
ì
í
î
ï
ð
ñ
ò
ó
ô
õ
ö
ø
œ
š
ù
ú
û
ü
ý
þ
ÿ
Α
Β
Γ
Δ
Ε
Ζ
Η
Θ
Ι
Κ
Λ
Μ
Ν
Ξ
Ο
Π
Ρ
Σ
Τ
Υ
Φ
Χ
Ψ
Ω
α
β
γ
δ
ε
ζ
η
θ
ι
κ
λ
μ
ν
ξ
ο
π
ρ
ς
σ
τ
υ
φ
χ
ψ
ω
ℵ
ϖ
ℜ
ϒ
℘
ℑ
←
↑
→
↓
↔
↵
⇐
⇑
⇒
⇓
⇔
∴
⊂
⊃
⊄
⊆
⊇
⊕
⊗
⊥
⋅
⌈
⌉
⌊
⌋
〈
〉
◊